John is on top of the building and Jack is down. If John throws ball at an angle of 60° and with initial velocity 20m/s at what height will the ball reach after 2s? Consider the motion along the y direction. Show your solution. (The height of the building was not given)


1
Expert's answer
2021-10-24T18:26:04-0400

The ball will move upward during


"t=\frac{v\sin{60}}{g}=1.77\ s"

The height above the building top is


"h=0.5gt^2=15.3\ m"

Then the ball falls during 2-1.77=0.23 s and covers a vertical distance of


"d=0.5gt^2_1=0.3\ m"

Thus, the ball will be at a height of


"H=h-d=15 \ m"

above the building.